# InkSight: Offline-to-Online Handwriting Conversion by Learning to Read and Write

### v2

@@ 4.1 Datasets
We train our models using two types of datasets: publicly available data and our in-house proprietary data. Below, we detail the public datasets and provide aggregated statistics for the in-house counterparts. Additional details are provided in Appendix B.
@@
As public OCR training data, we use RIMES (Augustin et al., 2006; Grosicki et al., 2009), HierText (Long et al., 2022, 2023), IMGUR5K (Krishnan et al., 2023), ICDAR’15 historical documents (Murdock et al., 2015), and IAM (Marti & Bunke, 1999). We crop out word-level images where possible, acquiring 295 000 samples with text in Latin script.
The in-house dataset consists of images of handwritten and printed text (67% and 33% respectively) written on a diverse set of backgrounds, with 95% of the labels in English.
@@
As public digital ink training data, we use VNOnDB (Nguyen et al., 2018), SCUT-Couch (Li et al., 2008), and DeepWriting (Aksan et al., 2018). For DeepWriting we use the available segmentation information to produce 3 datasets with character-level, word-level, and line-level croppings, and use all 3 for training. For VNOnDB we extract individual words to be used as training data. The public dataset size totals $\sim$2.7M samples.
The in-house multilingual dataset primarily comprises short snippets of text in Mandarin (37%) and Japanese (23%), with other languages contributing under 5%.
@@
Since there are no available diverse datasets of paired inks and images (as discussed in Section 2), we perform the evaluation on OCR data, and additionally do a small data collection to obtain paired samples.
To automatically assess our models, we evaluate the quality of derendering on the test splits of 3 OCR datasets: IAM (testset_f, $\sim$17.6k samples), IMGUR5K ($\sim$23.7k samples), and HierText. For HierText, which was not originally designed around handwritten samples, we apply the same filtering as to the OCR training data, and additionally only consider words marked as handwritten ($\sim$1.3k samples).
@@ 4.2 Models
We train 3 variants of our model: Small-i ($\sim$340M parameters, -i stands for in-house setup) with a ViT B/16 encoder pretrained on JFT-300M (Sun et al., 2017) paired with an mT5-base encoder-decoder; Small-p (-p stands for public setup), built with the same architecture as Small-i, but trained on public datasets and using a publicly available ImageNet-21k (Deng et al., 2009) pretrained ViT B/16 checkpoint; Large-i ($\sim$1B parameters) with a ViT L/16 encoder pretrained on JFT-300M, paired with an mT5-large encoder-decoder. All models employ a context length of 1024 for the output, and 128 for the input. Implementation details are provided in Appendix F, and a model card for Small-p in Appendix K.
@@ 4.3 Baseline Method
Despite pen trajectory recovery being a popular research direction in the last decades, available comparisons are limited due to the scarcity of open-sourced code, training data, and pre-trained model weights. Appendix C offers a detailed analysis of recent work and their reproducibility. We choose the General Virtual Sketching (GVS) Framework (Mo et al., 2021) as our baseline. Originally trained for sketch and photo vectorization, GVS offers a valuable reference point despite the domain discrepancy between sketches and handwritten notes. We found that providing the GVS line drawing vectorization model with binarized inputs yields best performance, we compare all 3 available models on original and binarized inputs in Appendix G.
@@ 4.4 Qualitative Evaluation
We compare the performance of our models and GVS on 3 public evaluation datasets (mentioned in Section 4.1), in Figure 5. Our models mostly produce results that accurately reflect the text content, disregarding semantically irrelevant background. They can also handle occlusions, highlighting the benefit of the learned reading prior, in contrast to GVS, which produces multiple duplicate strokes, and does not distinguish between background and foreground. Large-i is able to retain more details and accommodate more diverse image styles.
 Refer to caption Refer to caption
Figure 5: Comparison between performance of GVS, Small-i, Small-p, and Large-i on 3 public evaluation datasets (IAM, IMGUR5K, HierText from top to bottom), more visualizations in Section L.1. In every word, character colors transition from red to purple, following the rainbow sequence; within each stroke, the shade progresses from darker to lighter.
@@
We study performance of our models on out-of-domain samples on simple sketches. We use the Vanilla Derender inference mode to obtain the inks. In Figure 6, we observe that our models partly generalize to sketches, but the performance varies significantly depending on the sample and has noticeable artifacts such as missing details or over-focusing on a segment and over-tracing it. They, however, demonstrate robustness to complicated backgrounds (e.g. the lighting of the coffee sample).
@@ 4.5 Quantitative Comparison
To support the claims about the performance of our model, we conduct both a human evaluation and an automated evaluation that compare the similarity of our model output to the original image and to real digital inks. There are no established metrics and benchmarks for this problem yet.
@@ 4.5.1 Human evaluation
We performed a human evaluation of the quality of the derendered inks produced by the three variants of our model. We used the “golden” human traced data on the HierText dataset as the control group and the output of our model on these samples as the experiment group. Evaluators were shown the original image alongside a rendered digital ink, which was either model-generated or human-traced (unknown to the evaluators). They were asked to answer two questions: (1) Is the digital ink output a reasonable tracing of the input image? (Answers: Yes, it’s a good tracing; It’s an okay tracing, but has some small errors; It’s a bad tracing, has some major artifacts); (2) Could this digital ink output have been produced by a human? (Answers: Yes; No). We provide a more detailed description of the task and the instructions in Appendix J. The evaluation included 16 individuals familiar with digital ink, but not involved in this research. Each sample was evaluated by 3 raters and aggregated with majority voting (with inter-rater reliability $\kappa:$ (1) 0.46, (2) 0.44 for the respective questions).
We demonstrate the results of our evaluation campaign in Figure 7. We observe better performance (good or okay rating, more realistic samples) when both the data and model size are scaled up. The most common imprecisions that lead to choosing ”okay tracing” over ”good tracing” include: a small number of extra strokes (derendering irrelevant elements present in the image), missing details (e.g. punctuation, a dot over i, j), and unnecessary double strokes (see Appendix J for examples).
@@ 4.5.2 Automated Evaluation
Automated metrics (described below) are presented in Table 2 and most importantly the ordering using the automated metrics of our models matches the results of the human evaluation.
Table 2: Automated metrics comparison between three variants of our model, GVS, and the “golden” human traced data.
@@
F1
Acc.
F1
Acc.
F1
Acc.
Small-p
0.65
0.60
0.47
0.33
0.53
0.37
Small-i
0.65
0.59
0.51
0.33
0.61
0.44
Large-i
0.66
0.58
0.51
0.32
0.61
0.46
GVS
0.69
0.02
0.51
0.01
0.55
0.01
Human*
0.64
0.74
@@
To capture how similar the output of our model is to the input image semantically and geometrically, we compute the Character Level F1 score by following the protocol of the Robust Reading Challenge (Rob, 2023). This measure captures both geometry and semantics by counting the character bounding boxes found by an OCR model (in the original input image and in the output of the model rendered as an image) that both overlap significantly and match content-wise. To obtain those character bounding boxes, we use the publicly available API (Clo, 2024).
We see that our models perform similarly to GVS on simpler black-on-white IAM dataset, but outperforms it on HierText which has more diverse background. We also note that the OCR engine used is not perfect, and due to errors in the OCR model and its sensitivity to line width, the value of this metric is 0.64 for the “golden” human traced data on HierText. Therefore, it can only be used to meaningfully distinguish models as long as their F1 score is fairly low.
@@
To assess the semantic consistency and geometric properties of our generated digital inks, we compute recognition Exact Match Accuracy with the state-of-the-art online handwriting recognizer (Carbune et al., 2020) trained on real digital ink data.
With results shown in Table 2, our models show superior accuracy over GVS (on which the online handwriting recognizer is struggling due to unnatural stroke number, length and order). We also observed that while the three model variants perform similarly on the IAM and IMGUR5K datasets, the Large variant exhibits better performance compared to the Small variants on the HierText dataset which has greater text source diversity. This aligns with our human evaluation.
Table 3: Online handwriting recognition results on IAMOnDB test set for Small-i trained on real digital inks (IAMOnDB train set), derendered digital inks (from IAM train set) and their combination.
IAMOnDB IAM derendered IAMOnDB + IAM derendered
CER 6.1 7.8 4.6
To further confirm the similarity of real and derendered digital inks in terms of low-level feature statistics, we train an online handwriting recognizer on inks derendered from the IAM training set. We measure the performance on the testset_f of IAMOnDB (Liwicki & Bunke, 2005) and compare it with the model trained on IAMOnDB, with the commonly used Character Error Rate metric (CER). As shown in Table 2, model trained on derendered inks only has CER performance that is worse, but not far off from the one trained on real data. Furthermore, the combination of derendered IAM inks and real IAMOnDB inks allows to acquire a more diverse training set, resulting in significantly lower CER. We verify that this observation holds for Small-p and Large-i (see Appendix H). We attribute the relatively far from state-of-the-art CER of 6.1 to the small size of the training set. The details of the online handwriting recognizer we train are given in Appendix H.
@@ 4.6 Ablation Studies
This section examines the performance variation across inference tasks and ablates the impact of augmentation, data mixture, and training strategies on the Small-i model’s performance. Conclusions and numerical results of the ablation study are presented in Section 4.6 and summarized below. For more details, see Appendix D.
Table 4: Ablation studies on Small-i. The first row shows the original performance (with inference mode Derender with Text). Subsequent rows display results for the Vanilla Derender and Recognize and Derender tasks (Figure 3), followed by the impact of various design choices on performance.
@@
F1
Acc.
F1
Acc.
F1
Acc.
Small-i
0.66 ±$\pm 0.07$
0.59 ±$\pm 0.01$
0.51 ±$\pm 0.09$
0.33 ±$\pm 0.02$
0.61 ±$\pm 0.07$
0.45 ±$\pm 0.01$
Vanilla
$0.61\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.53\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.44\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.28\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.53\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.35\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
R+D
$0.62\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.57\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.46\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
0.32
$0.57\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.42\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
data aug*
$0.42\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.33\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.21\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.09\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.23\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.13\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
syn rec
$0.58\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.25\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.56\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.38\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
real rec
$0.64\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.55\,{\color[rgb]{0,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0}\pgfsys@color@rgb@stroke{0}{0.4}{0}\pgfsys@color@rgb@fill{0}{0.4}{0}\uparrow}$
$0.19\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.59\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.36\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
all rec
0.650.650.65\,0.65
$0.51\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
$0.55\,{\color[rgb]{0,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0}\pgfsys@color@rgb@stroke{0}{0.4}{0}\pgfsys@color@rgb@fill{0}{0.4}{0}\uparrow}$
$0.22\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
0.610.610.61\,0.61
$0.38\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\pgfsys@color@rgb@stroke{0.8}{0}{0}\pgfsys@color@rgb@fill{0.8}{0}{0}\downarrow}$
fr ViT
0.65 ±$\pm 0.13$
@@
0.43 ±$\pm 0.24$
0.31 ±$\pm 0.06$
0.59 ±$\pm 0.15$
@@
Inference task matters (Row 2-3).
We compare the performance between three types of inference tasks, we find Derender with Text is better than both Vanilla Derender and Recognize and Derender (R+D) on all three datasets, we attribute this difference to the fact that it generates more semantically consistent inks than the other two. Figure 8 shows a collection of derendered inks from Small-i where each input image has ambiguous or difficult-to-read characters. Inference with Vanilla Derender gives results that capture the overall geometry of the input image, but can lack textual understanding. Derender with Text gives results that are consistent with the text input to the model provided by OCR, while Recognize and Derender gives results that are consistent with the model’s own recognition. More examples in Section D.2.

### v4

@@ 4.1 Datasets
We train two different models using two types of datasets: one trained only on publicly available data and one trained on our in-house proprietary data. Below, we detail the public datasets and provide aggregated statistics for the in-house counterparts. Additional details are provided in Appendix C.
@@
As public OCR training data, we use RIMES (Augustin et al., 2006; Grosicki et al., 2009), HierText (Long et al., 2022; 2023), IMGUR5K (Krishnan et al., 2023), ICDAR’15 historical documents (Murdock et al., 2015), and IAM (Marti & Bunke, 1999). We crop out word-level images where possible, acquiring 295 000 samples with text in Latin script. The in-house dataset consists of images of handwritten and printed text (67% and 33% respectively) written on a diverse set of backgrounds, with 95% of the labels in English.
@@
As public digital ink training data, we use VNOnDB (Nguyen et al., 2018), SCUT-Couch (Li et al., 2008), and DeepWriting (Aksan et al., 2018). For DeepWriting we use the available segmentation information to produce 3 datasets with character-level, word-level, and line-level croppings, and use all 3 for training. For VNOnDB we extract individual words to be used as training data. The public dataset size totals $\sim$2.7M samples.The in-house multilingual dataset primarily comprises short snippets of text in Mandarin (37%) and Japanese (23%), while the remaining languages each contribute less than 5%.
@@
Since there are no available diverse datasets of paired inks and images (as discussed in Section 2), we perform the evaluation on OCR data and additionally do a small data collection to obtain paired samples. To automatically assess our models, we evaluate the quality of derendering on the test splits of 3 OCR datasets: IAM (testset_f, $\sim$17.6k samples), IMGUR5K ($\sim$23.7k samples), and HierText. For HierText, which was not originally designed around handwritten samples, we apply the same filtering as to the OCR training data, and additionally only consider words marked as handwritten ($\sim$1.3k samples).
@@ 4.2 Models
We train 3 variants of our model: Small-p ($\sim$340M parameters, -p stands for public setup) trained on publically available datasets with a ViT B/16 encoder pretrained on ImageNet-21k (Deng et al., 2009) paired with an mT5-base encoder-decoder; Small-i (-i stands for in-house setup), built with the same architecture as Small-p, but using a JFT-300M (Sun et al., 2017) pretrained ViT B/16 checkpoint; Large-i ($\sim$1B parameters) with a ViT L/16 encoder pretrained on JFT-300M, paired with an mT5-large encoder-decoder. All models employ a context length of 1024 for the output, and 128 for the input. Implementation details are provided in Appendix G, and a model card for Small-p in Appendix M.
@@ 4.3 Baseline Method
Despite pen trajectory recovery being a popular research direction in the last decades, available comparisons are limited due to the scarcity of open-sourced code, training data, and pre-trained model weights. Appendix D offers a detailed analysis of recent work and their reproducibility. We choose the General Virtual Sketching (GVS) Framework (Mo et al., 2021) as our baseline. Originally trained for sketch and photo vectorization, GVS offers a valuable reference point despite the domain discrepancy between sketches and handwritten notes. We found that providing the GVS line drawing vectorization model with binarized inputs yields best performance, we compare all 3 available models on original and binarized inputs in Appendix H.
@@ 4.4 Qualitative Evaluation
We compare the performance of our models and GVS on 3 public evaluation datasets (mentioned in Section 4.1) in Figures 5 and 6. Our models mostly produce results that accurately reflect the text content, disregarding semantically irrelevant background. They can also handle occlusions, highlighting the benefit of the learned reading prior, in contrast to GVS, which produces multiple duplicate strokes, and does not distinguish between background and foreground (which is reasonable, given it was trained for sketch vectorization). Large-i is able to retain more details and accommodate more diverse image styles.
Figure 5: Comparison between performance of GVS, Small-i, Small-p, and Large-i on IAM (top) and IMGUR5K (bottom), more visualizations in Section N.1.
Figure 6: Comparison between performance of GVS, Small-i, Small-p, and Large-i on HierText, which includes data with varying styles and occlusions.
@@
We additionally study the performance of our models on out-of-domain samples, such as simple sketches. For this, we use the Vanilla Derender inference mode to obtain the inks. In Figure 7, we observe that our models partly generalize to sketches, but the performance varies significantly depending on the sample and has noticeable artifacts such as missing details or over-focusing on a segment and over-tracing it. They, however, demonstrate robustness to complicated backgrounds (e.g. the lighting of the coffee sample).
@@ 4.5 Quantitative Comparison
To support the claims about the performance of our model, we conduct both human evaluation and automated evaluation that compare the similarity of our model output to the original image and to real digital inks. To the best of our knowledge, there are no established metrics and benchmarks for this problem yet.
@@ 4.5.1 Human evaluation
We performed a human evaluation of the quality of the derendered inks produced by the three variants of our model. We used the “golden” human traced data on the HierText dataset as the control group and the output of our model on these samples as the experiment group. Evaluators were shown the original image alongside a rendered digital ink, which was either model-generated or human-traced (unknown to the evaluators). They were asked to answer two questions:
  1. 1.
    Is the digital ink output a reasonable tracing of the input image? (Answers: Yes, it’s a good tracing; It’s an okay tracing, but has some small errors; It’s a bad tracing, has some major artifacts);
  2. 2.
    Could this digital ink output have been produced by a human? (Answers: Yes; No).
We provide a more detailed description of the task and the instructions in Appendix L. The evaluation included 16 individuals familiar with digital ink, but not involved in this research. Each sample was evaluated by 3 raters and aggregated with majority voting (with inter-rater reliability $\kappa:$ (1) 0.46, (2) 0.44 for the respective questions).
We demonstrate the results of our evaluation campaign in Figure 8. We observe better performance (good or okay rating, more realistic samples) when both the data and model size are scaled up. The most common imprecisions that lead to choosing “okay tracing” over “good tracing” include: a small number of extra strokes (e. g. derendering irrelevant elements present in the image), missing details (e.g. punctuation, a dot over i, j), and unnecessary double strokes (see Appendix L for examples).
@@ 4.5.2 Automated Evaluation
Automated metrics (described below) are presented in Table 2. Most importantly, the model ranking using the automated metrics matches the results of the human evaluation.
Table 2: Automated metrics comparison between three variants of our model, GVS, and the “golden” human traced data. We compare both Character Level F1 score of the rendered characters and Accuracy of recognizing the generated digital ink with state-of-the-art recognizers.
@@
F1 Acc. F1 Acc. F1 Acc.
Small-p 0.65 0.60 0.47 0.33 0.53 0.37
Small-i 0.65 0.59 0.51 0.33 0.61 0.44
Large-i 0.66 0.58 0.51 0.32 0.61 0.46
GVS 0.69 0.02 0.51 0.01 0.55 0.01
Human* 0.64 0.74
@@
To capture how similar the output of our model is to the input image semantically and geometrically, we compute the Character Level F1 score by following the protocol of the Robust Reading Challenge (Rob, 2023). This measure captures both geometry and semantics by counting the character bounding boxes found by an OCR model (in the original input image and in the output of the model rendered as an image) that both overlap significantly and match content-wise. To obtain those character bounding boxes, we use the publicly available API (Clo, 2024). We see that our models perform similarly to GVS on simpler black-on-white IAM dataset, but outperforms it on HierText which has more diverse background. We also note that the OCR engine used is not perfect, and due to errors in the OCR model and its sensitivity to line width, the value of this metric is 0.64 for the “golden” human traced data on HierText. Therefore, it can only be used to meaningfully distinguish models as long as their F1 score is fairly low.
@@
To assess the semantic consistency and geometric properties of our generated digital inks, we compute recognition Exact Match Accuracy with the state-of-the-art online handwriting recognizer trained on real digital ink data (Clo, 2024). As shown in Table 2, our models demonstrate superior accuracy compared to GVS (on which the online handwriting recognizer is failing completely due to unnatural stroke number, length and order). We also observed that while the three model variants perform similarly on the IAM and IMGUR5K datasets, the Large variant outperforms Small variants on the HierText dataset which has greater diversity, which aligns with our human evaluation.
To further validate the similarity between real and derendered digital inks, we train an online handwriting recognizer on inks derendered from the IAM training set. Testing on the testset_f of IAMOnDB (Liwicki & Bunke, 2005) and compare with the performance of model trained on IAMOnDB train set, in terms of Character Error Rate metric (CER). As shown in Figure 9, the model trained only on derendered inks achieves slightly worse CER performance compared to the model trained on real data. Furthermore, the combination of derendered IAM inks and real IAMOnDB inks allows to acquire a more diverse training set, resulting in significantly lower CER. This results holds for Small-p and Large-i as well (see Appendix I). While our CER trails state-of-the-art CER due to limited training data, our work demonstrates that offline handwriting datasets can effectively augment online recognizer training. The details of the online handwriting recognizer we train are given in Appendix I.
Figure 9: Online handwriting recognition results on IAMOnDB test set for Small-i trained on real digital inks (IAMOnDB train set), derendered digital inks (from IAM train set) and their combination.
@@ 4.6 Ablation Studies
This section examines the performance variation across inference tasks and ablates the impact of augmentation, data mixture, and training strategies on the Small-i model’s performance. Conclusions and numerical results of the ablation study are presented in Section 4.6 and summarized below.
Table 3: Ablation studies on Small-i. The first row shows the best performance setup (with inference mode Derender with Text). Subsequent rows display results for the Vanilla Derender and Recognize and Derender (R + D) tasks (Figure 3), followed by the impact of various design choices on performance.
@@
F1 Acc. F1 Acc. F1 Acc.
Small-i 0.66 ±$\pm 0.07$ 0.59 ±$\pm 0.01$ 0.51 ±$\pm 0.09$ 0.33 ±$\pm 0.02$ 0.61 ±$\pm 0.07$ 0.45 ±$\pm 0.01$
Vanilla $0.61\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.53\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.44\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.28\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.53\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.35\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
R+D $0.62\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.57\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.46\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ 0.32 $0.57\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.42\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
@@
data aug* $0.42\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.33\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.21\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.09\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.23\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.13\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
syn rec $0.58\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.25\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.56\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.38\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
real rec $0.64\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.50\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.55\,{\color[rgb]{0,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0}\uparrow}$ $0.19\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.59\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.36\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
all rec 0.650.650.65\,0.65 $0.51\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ $0.55\,{\color[rgb]{0,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0}\uparrow}$ $0.22\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$ 0.610.610.61\,0.61 $0.38\,{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}\downarrow}$
@@
For more details, see Appendix E.
@@
Inference task matters in difficult cases (Rows 2-3).
We compare the performance between three types of inference tasks, we find Derender with Text is slightly better than Recognize and Derender (R+D), which itself outperforms Vanilla Derender on all three datasets, we attribute this difference to the fact that it generates more semantically consistent. Figure 10 shows a collection of derendered inks from Small-i where each input image has ambiguous or difficult-to-read characters. We take the first column, DESiGNÉRS, as an example: Inference with Vanilla Derender captures the overall geometry of the input image but often lacks textual understanding (e.g., the letter E and ÉR are drawn as rough strokes rather than following the precise structure of the characters). Derender with Text produces results consistent with the text input to the model provided by OCR (upper case E and ER), while Recognize and Derender aligns with the model’s own recognition, resulting in lowercase e and er). More examples in Section E.2.